Twenty pounds of potatoes are divided between 2 families. One family receives 2 pounds more than the other. How many pounds did one of the families receive?
Answer by fractalier(6550)

Call what one family receives x.
The other family would then get 20 - x.
Then, since one quantity is two more than the other, we write
x + 2 = 20 - x
2x = 18
x = 9
x + 2 = 11
The families got 9 and 11 pounds.